A bag contains 3 red marbles, 6 blue marbles, and 1 white marble. the marbles are selected at random, one at a time, and ARE replaced. find the probability of....
A- blue and red
B- blue and not red

Answer by Alan3354(69443)

A bag contains 3 red marbles, 6 blue marbles, and 1 white marble. the marbles are selected at random, one at a time, and ARE replaced. find the probability of....
A- blue and red = (6/10)*(3/10) = 9/50
B- blue and not red = (6/10)*(7/10) = 21/50
------------